# Basis Number of Graphs Excluding Minors

### v2

@@ 3 Adhesions captured by a path system
Let us now prove that $\mathcal{B}$ indeed generates the cycle space of $G$. The arguments we present are standard, and similar to the ones used in [11] and [8, Section 5.3].
@@ 7.1.2 Facial walks
A central concept introduced by Mohar and Thomassen [21, Section 4] when working on graphs in surfaces is the notion of combinatorial embedding, which allows to describe topological embeddings of graphs on a surface in a purely combinatorial way, up to homeomorphism. We stress out that [21] focuses on cellular embeddings of graphs on surfaces, while our proof will require to consider general embeddings. However, the notions and results we will use here also hold in this settings, and we refer to [12] for a more complete and formal treatment of general embeddings.
@@ 8.1 Graphs with unbounded basis-number
The next result, even though not stated explicitly, follows directly from the proof of [26, Theorem 3].
@@
Lemma 8.1 implies in particular that any family of graphs with minimum degree 3 and unbounded girth has unbounded basis number. Such graph classes can be obtained using the probabilist method [7], or with some more explicit constructions [25] (see [23, Theorem 2.13] for a simple proof). Another way of obtaining explicit constructions, is to consider Ramanujan expanders, which moreover are Cayley graphs, and thus enjoy stronger algebraic properties. The first known constructions of Ramanujan expanders are due to Lubotzky, Phillips, Sarnak and independently Margulis [16, 18]. In particular, a characteristic of these constructions is that they produce graphs with girth $\Omega(\log(n))$. Next Theorem sums up some properties of such families, enabling to apply Lemma 8.1. The constructions from next theorem were provided by Morgenstern [22], and generalise the aforementioned ones to obtain infinite families of Ramanujan graphs which are $(q+1)$-regular for any arbitrary prime power $q$ (see also [4] for the case $q=2$).
@@
To conclude the proof, we set $\mathcal{G}_{1}\coloneqq\mathcal{H}_{2}$ and $\mathcal{G}_{2}\coloneqq\bigcup_{q}\mathcal{H}_{q}$, where the union is taken over all prime powers $q$. Clearly, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ have unbounded basis number. Moreover, a result of Godsil and Royle [9, Theorem 3.4.2] states that for every $d\geqslant 0$, every finite vertex-transitive $d$-regular is $\lceil\tfrac{2}{3}(d+1)\rceil$-connected. This implies in particular that $\mathcal{G}_{2}$ contains graphs of arbitrary large connectivity. ∎
@@ 8.2 Future directions
As pointed out in Corollary˜1.3, excluding minors characterises bounded basis number for subgraph closed graph classes. To go beyond this, it seems natural to consider graph classes that are closed only under taking induced subgraphs — or equivalently closed under induced minors, as edge contractions do not increase the basis number. A full characterisation of induced-minor closed classes with bounded basis number seems well beyond reach, but it would be interesting to find examples or establish properties of such classes. For instance, generalising Corollary˜1.3, one may ask whether basis number can be upper-bounded by a function of other classical graph parameters such as clique-width. Let us point out that such a bound cannot exist for the more general parameter twin-width, as subdividing all edges sufficiently many times decreases the twin-width of any graph to 4 [2] without affecting its basis number.
@@
Finally, basis number can be generalised to infinite graphs. Thomassen [28, Theorem 7.4] generalised MacLane’s planarity criterion, by proving that the infinite graphs having basis number at most $2$ are exactly those graphs admitting a vertex accumulation free planar embedding, that is, an embedding in $\mathbb{R}^{2}$ such that every point of $\mathbb{R}^{2}$ has an open neighbourhood containing only finitely many vertices. However, it is not clear whether in the infinite setting, arbitrary planar graphs, graphs embedded in a fixed surface, bounded treewidth graph, or graphs excluding a fixed minor have bounded basis number. As a side remark, observe that every Cayley graph of a finitely presented group has finite basis number. In particular, as planar and more generally minor-excluded finitely generated groups are finitely presented [5, 8], this question admits a positive answer when restricted to Cayley graphs (and more generally locally finite quasi-transitive graphs).
@@ Acknowledgements
We are deeply grateful to Babak Miraftab, Pat Morin and Yelena Yuditsky for very instructive exchanges, and for sharing with us their results once we learned that part of our respective works overlapped. Moreover, we thank Pat Morin for pointing to us a bound for Lemma 2.7 simplifying our original bound. We also thank Louis Esperet for some useful discussion, that lead in particular to a simplification of our initial proof of Corollary 7.6.

### v3

@@ 3 Adhesions captured by a path system
Let us now prove that $\mathcal{B}$ indeed generates the cycle space of $G$. The arguments we present are standard, and similar to the ones used in [15] and [12, Section 5.3].
@@ 7.1.2 Facial walks
A central concept introduced by Mohar and Thomassen [25, Section 4] when working on graphs in surfaces is the notion of combinatorial embedding, which allows to describe topological embeddings of graphs on a surface in a purely combinatorial way, up to homeomorphism. We stress out that [25] focuses on cellular embeddings of graphs on surfaces, while our proof will require to consider general embeddings. However, the notions and results we will use here also hold in this settings, and we refer to [16] for a more complete and formal treatment of general embeddings.
@@ 8.1 Graphs with unbounded basis-number
The next result, even though not stated explicitly, follows directly from the proof of [30, Theorem 3].
@@
Lemma 8.1 implies in particular that any family of graphs with minimum degree 3 and unbounded girth has unbounded basis number. Such graph classes can be obtained using the probabilist method [11], or with some more explicit constructions [29] (see [27, Theorem 2.13] for a simple proof). Another way of obtaining explicit constructions, is to consider Ramanujan expanders, which moreover are Cayley graphs, and thus enjoy stronger algebraic properties. The first known constructions of Ramanujan expanders are due to Lubotzky, Phillips, Sarnak and independently Margulis [20, 22]. In particular, a characteristic of these constructions is that they produce graphs with girth $\Omega(\log(n))$. Next Theorem sums up some properties of such families, enabling to apply Lemma 8.1. The constructions from next theorem were provided by Morgenstern [26], and generalise the aforementioned ones to obtain infinite families of Ramanujan graphs which are $(q+1)$-regular for any arbitrary prime power $q$ (see also [5] for the case $q=2$).
@@
To conclude the proof, we set $\mathcal{G}_{1}\coloneqq\mathcal{H}_{2}$ and $\mathcal{G}_{2}\coloneqq\bigcup_{q}\mathcal{H}_{q}$, where the union is taken over all prime powers $q$. Clearly, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ have unbounded basis number. Moreover, a result of Godsil and Royle [13, Theorem 3.4.2] states that for every $d\geqslant 0$, every finite vertex-transitive $d$-regular is $\lceil\tfrac{2}{3}(d+1)\rceil$-connected. This implies in particular that $\mathcal{G}_{2}$ contains graphs of arbitrary large connectivity. ∎
@@ 8.2 Future directions
In terms of lower bounds, we note that for every $t\in\mathbb{N}$, there exists a graph with basis number $\Omega(\log(t))$ that excludes $K_{t}$ as a minor. This can be seen using for example [6, Proposition 2.2], which states that there exists a constant $c\in\mathbb{R}$ such that for infinitely many $t\in\mathbb{N}$, there exists a cubic graph $G_{t}$ with girth at least $\tfrac{8}{3}\log(t)-c$ that excludes $K_{t}$ as a minor. In particular, Lemma˜8.1 implies that such graphs have basis number $\Omega(\log(t))$. Note that we cannot hope for any better lower bound using this argument, as the authors from [6] showed the existence of a constant $c\in\mathbb{R}$ such that such that for all $t\in\mathbb{N}$, every graph with girth at least $6\log(t)+3\log\log(t)+c$ and minimum degree at least $3$ admits $K_{t}$ as a minor.
Hereditary classes
As pointed out in Corollary˜1.3, the property of being proper minor closed characterises basis number boundedness for monotone graph classes. To go beyond this, it seems natural to consider graph classes that are hereditary, that is, closed only under taking induced subgraphs — or equivalently closed under taking induced minors, as edge contractions do not increase the basis number. A full characterisation of induced-minor closed classes with bounded basis number seems well beyond reach, but we believe that it would be interesting to find examples or establish properties of such classes. For instance, generalising Corollary˜1.3, one may ask whether basis number can be upper-bounded by a function of other classical graph parameters such as clique-width. Let us point out that such a bound cannot exist for the more general parameter twin-width, as subdividing all edges sufficiently many times decreases the twin-width of any graph to 4 [2] without affecting its basis number.
@@
Infinite graphs
Finally, cycle bases and basis number can be generalised to infinite graphs. In fact, two different definitions have been considered in the litterature. The first one is more restrictive and defines the cycle space of a graph as its set of finite even subgraphs, equipped with $\oplus$. With respect to this notion of cycle space, Thomassen [32, Theorem 7.4] generalised MacLane’s planarity criterion, by proving that the infinite graphs having basis number at most $2$ are exactly those graphs admitting a vertex accumulation free planar embedding, that is, an embedding in $\mathbb{R}^{2}$ such that every point of $\mathbb{R}^{2}$ has an open neighbourhood containing only finitely many vertices. A natural question is whether planar graphs, and more generally graphs embedded in a fixed surface of bounded genus, bounded treewidth graphs, or graphs excluding a fixed minor have bounded (or even finite) basis number. Nathan Bowler (private communication) showed us a proof that with respect to this notion, infinite planar graphs have basis number at most $4$. As a side remark, observe that every Cayley graph of a finitely presented group has finite basis number. In particular, as planar and more generally minor-excluded finitely generated groups are finitely presented [9, 12], minor-excluded Cayley graphs (and more generally locally finite quasi-transitive graphs) have finite basis number.
A second possible definition of cycle space, called topological cycle space consists in considering all subgraphs generated by infinite cycles. We refer to [8, Section 8.7] and [7] for a comprehensive exposition on the topic, and to [4] for a proof of a generalisation of MacLane’s planarity criterion in this case. Again, basis number can be defined indentically with respect to the topological cycle space, and it makes sense to ask whether all aforementioned classes of graphs also have bounded basis number in this context.
@@ Acknowledgements
We are deeply grateful to Babak Miraftab, Pat Morin and Yelena Yuditsky for very instructive exchanges, and for sharing with us their results once we learned that part of our respective works overlapped. Moreover, we thank Pat Morin for pointing to us a bound for Lemma 2.7 simplifying our original bound. We also Louis Esperet for some useful discussion, that lead in particular to a simplification of our initial proof of Corollary 7.6, and Nathan Bowler for showing us a proof that infinite planar graphs have basis number at most $4$.